UT Mar 01 2021

Night Partition

  • Night split with LBTI - night center 00:32 - handover 00:27 to LBTO Eng LUCI-SOUL-AO
  • Moon is up effectively all night (94.6%)
  • 1st half LBTI
  • 2nd half LUCI-1 SOUL-AO

Night Plan Eng Time

AO Engineering Time - highest priority - (LBTO AO and Arcetri)
  • LUCI-SOUL AO tasks - test 6 different PAs per object
  • LUCI-SOUL Arcetri tests
    • closed loop on AO ref stars (on-axis) to collect SR and telmetry
    • one or more stars in the following categories
      • 14.5 < Rmag < 15.5 Fast Bin 3
      • 15.8 < Rmag < 16.5 Slow Bin 3
      • 16.51 < Rmag < 17.0 Fast Bin 4
      • 17.0 < Rmag < 18.0 Slow Bin 4
      • Rmag > 18.0 (as far as we can go....)
  • LUCI-SOUL ESM tasks - 1hr - only if no other full AO
    • Hoag's Object - on-axis AO Ref star (galaxy core as AO Ref star R=13.44 predicted WFSmag =14.7)
  • if cannot switch to LUCI due to M3 SX issue
    • MODS Field Correction check for IT 8315
      • 1-1.5 hours in seeing better than 1.4 arcsec, clouds don't matter much.
      • priority=medium (problem may have been with us for a year)
      • run IDL task field_collect (same as measuring field aberrations: ComTools#Field_Data_Collection)
      • do LDG and RDG in parallel (two IDL windows)
      • Use 6-8 stars (some fields only have 6 stars; start the list a second time, if you are under an hour)
      • Stone_G is well placed at RA=10:17, and other Stone / BS fields are available if earlier or later.
    • mid-z ULIRG FF0937+3857 - 2hrs MODS (Rothberg)
    • Local Mergers - 1hr per target - MODS (Rothberg)
    • LBC Solar System Objects - 1-2hrs (Veillet)
  • Poor Weather Backup
    • ESM data collection (Bino-LUCIs 1hr per target)
      • M67 visible from handover until 03hr
      • M92 visible from 2am through twilight
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/ESM_Testing
    • MODS Local Mergers (Bino-MODS 90min or longer per target based on conditions) - Rothberg
      • Script Location: /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/Local_Mergers/scripts
    • ICRF3 (bino-LUCIs) - Rothberg (if already switched to LUCI)
      • ESM has priority over ICRF3 - if similar conditions collect ESM data
      • /home/lbto/Service_Observing/2021A_D_E/Feb_Mar_2021/Rothberg_Science/ICRF3


LBTI first half, but starting up MODS & LUCI in the afternoon. Waking MODS, simSnap & starting LUCI, taking some darks and verifying fieldstop alignment. There are clouds moving in just before sunset anf moderate to high winds.

7:27 Reconfiguing for LUCI. And the M3 failed to reach position. Disucussions about how to proceed.

Calling Mike Gardiner but no answer.

7:49 We are reconfiguring for MODS.

8:06 David is performing pointing and collimation

8:17 Preset to Feige 34. And we have winds of 23m/s. We are closing. That was fun

10:17 LUCI1 alarming. LUCI1 status looks fine. No obvious cause. Apparently it did this earlier and rick investigated but found no cause then. We will continue to monitor LUCI1... Winds are dropping

10:50 Winds have dropped. David is opening

11:02 Rick is going to check on LUCI1

11:18 Rick is done. Compressor on LUCI1 had tripped. Slewing to BD33. On second thought, nope. Must have caught a gust, the shell ripped. And we are seeing winds over the closing again.

11:51. Winds continue to gust regularly over 22m/s. Putting MODS to bed.

-- BarryRothberg - 28 Feb 2021
Topic revision: r5 - 01 Mar 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback